OpenStudy (anonymous):

use the foil method to find the product (4x^7-3)(x^8-1)

OpenStudy (anonymous):

first: 4x^7 * x^8 = 4x^15 outer: 4x^7 * (-1) = -4x^7 inner: -3 * x^8 = -3x^8 last: -3 * (-1) = 3